OpenStudy (anonymous):

Blank ( X plus or minus Blank )= 6X plus or minus Blank, Answer: X = 4 How to complete the blanks? Thanks

OpenStudy (anonymous):

y(x-y)=6x-y; y(4-y)=24-y 4y-y^2=24-y y^2-5y+24=0 (y +3)(y -8)=0 so y either -3 or 8 y(x+y)=6x-y; y(4+y)=24-y 4y+y^2=24-y y^2+5y-24=0 (y -3)(y +8)=0 so y either +3 or- 8 i am not sure about Q

OpenStudy (anonymous):

so the blank +-3 and +-8